OpenStudy (anonymous):

find a coterminal angle between 0 and 2 pi for 11 pi/3

OpenStudy (lgbasallote):

is that possible? dont you add or subtract 2 pi to get coterminal angles?

OpenStudy (asnaseer):

\[\frac{11\pi}{3}=(3\frac{2}{3})\pi\]so just subtract \(2\pi\) from this to get the coterminal angle.